please help me under stand how to do this
4x^4-11x^2+6=0
thank you

Answer by mananth(16946)

4x^4-11x^2+6=0
a=4,b=-11 c=6
ac=24
split b such that the sum of split terms = -11 and product of the split terms = -24
-8 & -3
-8*-3=24
-8+(-3)=-11
4x^4-8x^2-3x^2+6
4x^2(x^2-2)-3(x^2-6)
(x^2-2)(4x^2-3)